OpenStudy (anonymous):

Evaluate. 9!/6!4! 14 21 42

OpenStudy (ash2326):

\[\frac{9!}{6!4!}\] We know that \[n!=n.(n-1)!\] \[\frac{9.8.7.6!}{6!4!}\] Can you solve now?

OpenStudy (lgbasallote):

\[\Large \frac{(9)(8)(7)(6)(5)(4)(3)(2)(1)}{(6)(5)(4)(3)(2)(1) \times (4)(3)(2)(1)}\] cancel all that an be

OpenStudy (lgbasallote):

can be*

OpenStudy (anonymous):

21 ;) thanks